OpenStudy (anonymous):

A 10 kg lead brick falls from a height of 2.0m. a) Find its momentum as it reaches the ground. b) What impulse is needed to bring the brick to rest? c) The brick falls onto a carpet, 1.0 cm thick. Assuming the force stopping it is constant, find the average force the carpet exerts on the brick. d) If the brick falls onto a 5.0 cm foam rubber pad, what constant force is needed to bring it to rest? Please help, I have tried.. and so far I got a) -40 kg x m/s ..Not sure if even that is right

OpenStudy (souvik):

first find the velocity of the brick as it reaches the ground..

OpenStudy (souvik):

@brightdiwa

OpenStudy (souvik):

\[v^2=u^2+2gh\] u=initial velocity=0 m/s v=final velocity g=10 m/s^2 h=distance=2.0 m \[v=2\sqrt{10}\] m/s so the momentum is mass*velocity=\(20\sqrt{10}\)

OpenStudy (anonymous):

where did u get that value for g?

OpenStudy (unklerhaukus):

\[g\approx9.8[\text {m/s}^2]\sim10[\text {m/s}^2]\]

OpenStudy (souvik):

you can use g=9.8 m/s^2...

OpenStudy (anonymous):

A 60 kg dancer leaps 0.32m high. With what momentum does the dancer reach the ground? What impulse is needed to make a stop? As the dancer lands, the knees bend, lengthing the stopping time to 0.050 seconds. FInd the average force exerted on the body. Compare the stopping force to the performer's weight. I need help on this question Any help and explanation is appreciated